Find the slope if it exists
x+8y=16
need to know how to solve this problem and how to either write it out in integer or simplified fraction

Answer by mananth(16946)

x+8y=16
you have to rewrite the equation in the form y = mx+b
where m is the slope and b the y intercept.
x+8y=16
add -x
x-x+8y=-x+16
8y= -x+16
/8
y = -x/8 +2
the slope is -(1/8)
..
m.ananth@hotmail.ca